Cashews sell for $4.00 per pound, Peanuts for $1.50 per pound. If you mix 60 lbs of peanuts in a cashew/peanut mix, how many pounds of cashews do you need to add to make a mix that costs $2.50 per pound.
Answer by ankor@dixie-net.com(22740)

: Cashews sell for $4.00 per pound, Peanuts for $1.50 per pound. If you mix 60 lbs of peanuts in a cashew/peanut mix, how many pounds of cashews do you need to add to make a mix that costs $2.50 per
:
Let x = lb of cashews required
Then
The final amt = (x+60)
:
A nutty equation:
4x + 1.5(60) = 2.5(x+60)
:
4x + 90 = 2.5x + 150
:
4x - 2.5x = 150 - 90
:
1.5x = 60
:
x = 60/1.5
:
x = 40 lb of cashews required
:
:
Check solution in original equation:
4(40) + 1.5(60) = 2.5(40+60)
160 + 90 = 2.5(100) proves our solution
:
:
Not that hard, right?